OpenStudy (hitaro9):

Does the Series ke^-k^2 converge or diverge?

OpenStudy (hitaro9):

\[\sum_{k=1}^{infinity}k e^{-k ^{2}}\]

OpenStudy (anonymous):

$$\large \sum_{k=1}^{\infty}k e^{-k ^{2}}= $$ notice that is close to the derivative of \( \large e^{-x^2} \)

OpenStudy (hitaro9):

Oh, right, there's a thing in the book about an integral test. We haven't discussed it in class yet, but, I'm guessing that's what this is?

OpenStudy (anonymous):

sort of, I was thinking of this http://prntscr.com/8j57eg

OpenStudy (hitaro9):

That looks significantly more complicated than what we've been doing thus far.

OpenStudy (anonymous):

but now that you mention it, you can use integral test

OpenStudy (hitaro9):

Okay then, that's good. I'm reading up on it now, might ask some questions

OpenStudy (irishboy123):

applied the ratio test yet?

OpenStudy (hitaro9):

Ratio test didn't seem to work, unless I made some algebraic errors

OpenStudy (irishboy123):

\[\large \frac{(k+1) e^{-(k+1) ^{2} }}{k e^{-k ^{2}}}=\]

OpenStudy (hitaro9):

Right, then I ended up with some mess with the exponents that didn't seem to work out nicely

OpenStudy (irishboy123):

\[\large \frac{k+1}{k} e^{-(2k+1)}\]

OpenStudy (hitaro9):

To save you the work of typing out a whole bunch of equations, if there's some way to do it with the ratios, I should be able to handle it if you can help me through the algebra

OpenStudy (irishboy123):

right so far?

OpenStudy (hitaro9):

Umm one moment. That's not what I did so that's probably where I went wrong

OpenStudy (irishboy123):

expand the top exponent and subtract the bottom one

OpenStudy (hitaro9):

Right that makes sense. I dunno why I didn't see that.

OpenStudy (anonymous):

because the e^(-2k+1) is in the denominator. nice work irishboy